OpenStudy (anonymous):

whats the derivative of 4lnx?

OpenStudy (anonymous):

4/x

OpenStudy (anonymous):

derivative of lnx = 1/x so the c will not change so it is 4/x product rule: (4)'(lnx) + 4(lnx)' = 0 + 4(1/x) = 4/x

jimthompson5910 (jim_thompson5910):

alternatively, you can pull out the constant 4 and take the derivative of ln(x) to get 1/x so you would get 4*(1/x) = 4/x